OpenStudy (anonymous):

Partial Derivatives: u= xy(sin^-1(yz)) find the first partial derivative with respect to x For the product rule how do I derive sin^-1(yz) in respect to x???

OpenStudy (kainui):

y and z are independent of x correct? Then this is simply a constant.

OpenStudy (jhannybean):

Use the product rule here. \(\ f'g + g'f\)

OpenStudy (anonymous):

Yes. So, it would be 0. @Kainui

OpenStudy (kainui):

Well if you look at it, really y(sin^-1(yz)) is all the coefficient on x. So you don't even need to use a product rule here.

OpenStudy (jhannybean):

\[\large U_x = (xy)'(sin^{yz}) + (sin^{yz})'(xy)\]

OpenStudy (jhannybean):

Hmm I guess. I use the product rule as a safety net, just so I know when I take the derivative with respect to x I can clearly see what cancels out, and whatnot.

OpenStudy (anonymous):

So, the answer would be y(sin^-1(yz)) @Kainui

OpenStudy (jhannybean):

And darn it, I wrote the \(\sin^{-1} (yz)\) a bit wrong <_<

OpenStudy (kainui):

Suppose you try to use the product rule, you'll get: \[\LARGE u=xy \sin^{-1}(yz)\] \[\Large \frac{\partial u}{\partial x}=\frac{d}{dx}(x)*y \sin^{-1}(yz)+x \frac{d}{dx}(y \sin^{-1}(yz))\] You get: \[\Large \frac{\partial u}{\partial x}=(1)*y \sin^{-1}(yz)+x* (0)\] which is the correct answer still.

OpenStudy (anonymous):

@Kainui Oh ok. I get it!! Thank you!

OpenStudy (kainui):

Cool, glad I could help! =P